COMMUTATIVE IDEAL THEORY WITHOUT FINITENESS CONDITIONS: COMPLETELY IRREDUCIBLE IDEALS

LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

Abstract. An ideal of a ring is completely irreducible if it is not the intersection of any set of proper overideals. We investigate the structure of completely irrreducible ideals in a commutative ring without finiteness conditions. It is known that every ideal of a ring is an intersection of completely irreducible ideals. We characterize in several ways those ideals that admit a representation as an irredundant intersection of completely irreducible ideals, and we study the question of uniqueness of such representations. We characterize those commutative rings in which every ideal is an irredundant intersection of completely irreducible ideals.

Introduction

Let R denote throughout a commutative ring with 1. An ideal of R is called irreducible if it is not the intersection of two proper overideals; it is called completely irreducible if it is not the intersection of any set of proper overideals. Our goal in this paper is to examine the structure of completely irreducible ideals of a commutative ring on which there are imposed no finiteness conditions. Other recent papers that address the structure and ideal theory of rings without finiteness conditions include [3], [4], [8], [10], [14], [15], [16], [19], [25], [26].

A proper ideal A of R is completely irreducible if and only if there is an element x ∈ R such that A is maximal with respect to not containing x. Indeed, the condition is clearly sufficient for A to be completely irreducible since x is in the intersection of the proper overideals of A. On the other hand, if A is completely irreducible and x is an element that is not in A but is in the intersection of the proper overideals of A, then A is maximal with respect to not containing x. Evidently, maximal ideals are completely irreducible. If R is a domain (not a field), then the zero ideal of R is prime and irreducible, but it is not completely irreducible (it is the intersection of all nonzero ideals). More generally, a prime ideal is always

Date: July 26, 2004.

1991 Mathematics Subject Classification. Primary 13A15, 13F05.

Key words and phrases. irreducible ideal, completely irreducible ideal, irredundant intersection, arithmetical ring .

2 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

irreducible and is completely irreducible if and only if it is a maximal ideal. In Theorem 1.3 we characterize completely irreducible ideals in various ways. We deduce in Corollary 1.5 that an irreducible ideal of a Noetherian ring is completely irreducible if and only if it is primary for a maximal ideal.

A great deal is known about the structure of the irreducible ideals of a Noetherian ring. Indeed, the first decomposition theorem established by Emmy Noether [27] states that each ideal of a Noetherian ring admits a representation as an irredundant intersection of finitely many irreducible ideals; moreover, the number of components appearing in such a representation is an invariant, and each such representation is reduced (i.e., no ideal in the representation can be replaced by a strictly larger ideal to obtain the same intersection). Another result due to Noether [27] is that a proper irreducible ideal of a Noetherian ring is a primary ideal. In a ring without finiteness assumptions there may exist proper irreducible ideals that are not primary. Fuchs in [12] introduced the concept of a primal ideal, where a proper ideal A of R is said to be primal if the zero-divisors in R/A form an ideal. The ideal of zero-divisors is then necessarily of the form P/A where P is a prime ideal of R called the adjoint prime of A. We also say that A is P -primal. In Fuchs [12] it is shown that proper irreducible ideals are primal. More can be said about completely irreducible ideals. We observe in Theorem 1.3 several equivalences to a proper ideal C of R being completely irreducible among which is that C is irreducible and R/C contains a simple submodule.

A ring in which every irreducible ideal is completely irreducible is zero-dimensional. Corollary 1.5 states that if dim R = 0 and each primary ideal of R contains a power of its radical, then every irreducible ideal of R is completely irreducible. In Section 2 we address the question: Under what conditions is an ideal representable as an irredundant intersection of completely irreducible ideals? We consider in Section 3 the question of uniqueness of representation of the ideal A as an irredundant intersection of completely irreducible ideals. In Section 4 we characterize the rings in which every ideal is an irredundant intersection of completely irreducible ideals. We prove in Theorem 4.2 that every ideal of the ring R is an irredundant intersection of completely irreducible ideals exactly if the ring is semi-artinian, where a ring R is said to be semi-artinian if every nonzero R-module contains a simple R-module.

Proposition 1.4 and Corollary 1.5 characterize the completely irreducible ideals of a Noetherian ring. In Section 5 we give an explicit description of the completely

COMPLETELY IRREDUCIBLE IDEALS 3

irreducible ideals of an arithmetical ring, where the ring R is arithmetical if for every maximal ideal M the ideals of the localization R_{M} are linearly ordered with respect to inclusion. An arithmetical integral domain is a Prüfer domain. For a prime ideal P of a Prüfer domain, Fuchs and Mosteig prove in [17, Lemma 4.3] that the P -primal ideals form a semigroup under ideal multiplication. We generalize this result in Theorem 5.6. For a prime ideal P of an arithmetical ring, we show that the regular P -primal ideals form a semigroup under ideal multiplication, where an ideal is regular if it contains a nonzerodivisor. Theorem 5.8 states that if M is a maximal ideal of an arithmetical ring R, then the set F of completely irreducible regular ideals of R with adjoint maximal ideal M is closed under ideal-theoretic multiplication, and F with this multiplication is a totally ordered cancellative semigroup.

A good reference for our terminology and notation is [18]. For ideals I, J of the ring R, the residual I : J is defined as usual by

I : J = {x ∈ R : xJ ⊆ I}.

For an ideal A and for a prime ideal P of R, we use the notation

A_{(P)} = {x ∈ R : sx ∈ A for some s ∈ R \ P } = ⋃_{s∈R\P} A : s

to denote the isolated P -component (isoliertes Komponentenideal) of A in the sense of Krull [24, page 16]. Notice that x ∈ A_{(P)} if and only if A : x ⊈ P. If R is a domain, then A_{(P)} = AR_{P} ∩ R, where R_{P} denotes the localization of R at P.

Two different concepts of associated primes of a proper ideal A of the ring R are useful for us. One of these was introduced by Krull [23, page 742], and following [22] we call a prime ideal P of R a Krull associated prime of A if for every x ∈ P, there exists y ∈ R such that x ∈ A : y ⊆ P. The prime ideal P is said to be a strong Bourbaki associated prime of A if P = A : x for some x ∈ R.

1. Irreducible and Completely Irreducible Ideals

A ring is called subdirectly irreducible if in any of its representations as a subdirect product of rings, one of the projections to a component is an isomorphism. It is straightforward to see:

Lemma 1.1. A proper ideal C of R is completely irreducible if and only if the factor ring R/C is subdirectly irreducible, or equivalently if and only if the R-module R/C has a simple essential socle. □

4 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

Thus if C is completely irreducible, then R/C contains a minimal nonzero ideal C^{∗}/C: the intersection of all nonzero ideals of R/C; this is then the essential socle of R/C. We shall call C^{∗} the cover of C. Clearly, C^{∗}/C is a simple R-module, so C^{∗}/C ≅ R/M for a maximal ideal M of R. (If C is only assumed to be irreducible, then we can only claim that R/C is a uniform R-module, that is, the intersection of any two nonzero submodules of R/C is not zero.)

Proposition 1.2. A completely irreducible proper ideal C of the ring R is a primal ideal whose adjoint prime is the maximal ideal M of R for which C^{∗}/C ≅ R/M. Furthermore, M is a strong Bourbaki associated prime of C.

Proof. Let x ∈ R be a representative of any coset of C generating C^{∗}/C ≅ R/M. Then x ∉ C and xM ⊆ C. Hence M ⊆ C : x ⊂ R (proper containment); thus we necessarily have M = C : x, so M is a strong Bourbaki associated prime of C. Evidently, M has to be the adjoint prime of the primal ideal C. □

Completely irreducible ideals can be characterized in various ways as we demonstrate in Theorem 1.3.

Theorem 1.3. For a proper ideal C of R, the following conditions are equivalent. (i) C is completely irreducible; (ii) the factor module R/C is an essential extension of a simple module; (iii) C is an irreducible ideal and R/C contains a simple R-submodule; (iv) C is an irreducible ideal and C is properly contained in C : M for some

maximal ideal M of R;

(v) C is irreducible with adjoint prime a maximal ideal M of R, and M = C : x for some x ∈ R \ C;

(vi) C = C_{(M)} for some maximal ideal M of R and CR_{M} is a completely irreducible ideal of R_{M}.

Proof. The equivalence of (i) and (ii) and of (iii) and (iv) is obvious, and so is the implication (i) ⇒ (iv). To prove (iii) ⇒ (i), observe that a simple submodule of a uniform module is an essential socle of the module.

By Proposition 1.2, (i) ⇒ (v) is clear. On the other hand, Condition (v) implies that x + C generates a simple R-submodule in R/C. By the uniformity of R/C this simple submodule is an essential socle of R/C, and (i) holds.

The equivalence of (v) and (vi) follows from the following two observations: (a) if M is a maximal ideal containing C, then (C : x)R_{M} = CR_{M} :_{R_{M}} x for any

COMPLETELY IRREDUCIBLE IDEALS 5

x ∈ R; and (b) C_{(M)} is irreducible if and only if CR_{M} is an irreducible ideal of R_{M} (see Remark 1.6 of [14]). □

A ring is Laskerian if every ideal has a finite primary decomposition.

Proposition 1.4. If every proper ideal of the ring R is an intersection of primary ideals (possibly an infinite intersection), then every completely irreducible proper ideal of R is primary for a maximal ideal. In particular, if R is a Laskerian ring (or a Noetherian ring), then every completely irreducible proper ideal of R is primary for a maximal ideal.

Proof. Let C be a completely irreducible ideal of R with adjoint prime M and let x ∈ R be a representative of any coset of C generating C^{∗}/C ≅ R/M. Then x ∉ C and xM ⊆ C. Since C is an intersection of primary ideals, there exists a primary ideal Q of R such that C ⊆ Q and x ∉ Q. Since C is completely irreducible, it follows that C = Q. Therefore C is a primary ideal that is M-primal, so C is M-primary. □

Corollary 1.5. Let M be a maximal ideal of the ring R and let C be an irreducible M-primary ideal. If M^{n} ⊆ C for some positive integer n, then C is completely irreducible. Thus if M is finitely generated, then every irreducible M-primary ideal is completely irreducible. Therefore an irreducible proper ideal of a Noetherian ring is completely irreducible if and only if it is primary for a maximal ideal.

Proof. If M^{n} ⊆ C for some positive integer n, then C ⊂ (C : M). Hence by Theorem 1.3(iv), C is completely irreducible. The last statement now follows from Proposition 1.4. □

Remark 1.6. We are interested in describing the rings in which every irreducible ideal is completely irreducible. Since prime ideals are irreducible and a prime ideal is completely irreducible if and only if it is maximal, the condition that every irreducible ideal in R is completely irreducible implies that all prime ideals of R are maximal and dim R = 0. If R is a reduced zero-dimensional ring, then R_{P} is a field for each P ∈ Spec R and every primal ideal of R is maximal. Therefore every irreducible ideal of a reduced zero-dimensional ring is completely irreducible.

Recall that a ring R is semi-artinian if every nonzero R-module contains a simple R-module. If R is semi-artinian, then all irreducible ideals of R are completely irreducible (see also Lemma 2.4 in Dilworth-Crawley [7]). The existence of a reduced

6 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

zero-dimensional ring R having no principal maximal ideals shows that a ring in which all irreducible ideals are completely irreducible need not be semi-artinian. In Example 1.7 we give two specific ways to obtain an example of such a ring R.

Example 1.7. (1) Let L denote the algebraic closure of the field ℚ of rational numbers and let D be the integral closure of the ring ℤ of integers in the field L. Fix a prime integer p and let R = D/\sqrt{pD}. Then R is a reduced zero-dimensional ring having the property that no maximal ideal of R is finitely generated.

(2) Let ℚ^{ω} denote the product of countably infinitely many copies of the field ℚ of rational numbers and let I denote the direct sum ideal of ℚ^{ω}. Then R = ℚ^{ω} /I is a reduced zero-dimensional ring in which no maximal ideal is finitely generated.

Remark 1.8. Let M be a maximal ideal of the ring R. By Theorem 1.3, an irreducible M-primary ideal A is completely irreducible if and only if A ⊂ (A : M). If (R, M) is a rank-one nondiscrete valuation domain and A = xR is a principal M-primary ideal, then A is irreducible, but not completely irreducible. (Since the value group G of R is a dense subset of ℝ, the value of x is the limit of smaller elements of G.) On the other hand, if dim R = 0 and if each primary ideal of R contains a power of its radical, then Corollary 1.5 implies that all irreducible ideals of R are completely irreducible.

Let A ⊆ C be ideals of the ring R. If C is completely irreducible, we call C a relevant completely irreducible divisor of A if A has a decomposition as the intersection of completely irreducible ideals in which C is relevant (i.e. it cannot be omitted). If A admits an irredundant decomposition with completely irreducible ideals, then all the ideals in this decomposition are relevant. On the other hand, a prime ideal that is not a maximal ideal has no relevant completely irreducible divisors. In Proposition 1.9 we characterize the relevant completely irreducible divisors of an ideal.

Proposition 1.9. A completely irreducible ideal C containing the ideal A is a relevant completely irreducible divisor of A if and only if the submodule C/A of R/A is not essential.

Proof. First suppose C is a relevant completely irreducible divisor of A, and A = C ∩ ⋂_{i∈I} C_{i} is a decomposition with completely irreducible ideals C_{i}, where C

COMPLETELY IRREDUCIBLE IDEALS 7

cannot be omitted. Then there exist an element x ∈ ⋂_{i∈I} C_{i} \ C and an r ∈ R such that rx ∈ C^{∗} \ C. Let u = rx and let \overline{u} denote the image of u in R/A. Then \overline{u} ∉ C/A, while if M denotes the maximal ideal of R such that C^{∗}/C ≅ R/M, then Mu ⊆ C ∩ ⋂_{i∈I} C_{i} = A implies R\overline{u} is a simple submodule of R/A. Hence the submodule C/A of R/A is not essential.

Conversely, assume C/A is not essential in R/A, i.e. there exists x ∈ R \ A such that (Rx + A) ∩ C = A. Write the ideal (Rx + A) as an intersection of completely irreducible ideals C_{i} (i ∈ I). Then A = C ∩ ⋂_{i∈I} C_{i}, where C is relevant. □

Corollary 1.10. For every completely irreducible ideal C containing the ideal A, the cover C^{⋆} satisfies: C^{⋆}/A is an essential submodule in R/A.

Proof. If C is not a relevant completely irreducible divisor of A, then C/A is an essential submodule of R/A, so C^{∗}/A is essential in this case. On the other hand, if C is relevant, then for any x ∈ R \ C, there exists r ∈ R such that rx ∈ C^{∗} \ C. □

Corollary 1.11. Let A be a proper ideal of the ring R. There exists a relevant completely irreducible divisor of A if and only if the socle Soc(R/A) ≠ 0.

Proof. Assume that C is a relevant completely irreducible divisor of A. If A = C, then Soc(R/A) = Soc(R/C) = C^{∗}/C ≠ 0. If A ⊂ C, then A = C ∩ B, where A ⊂ B. Let b ∈ B \ A. Then b ∉ C, so there exists r ∈ R such that rb ∈ C^{∗} \ C. It follows that Mrb ⊆ C ∩ B = A, where M is the adjoint prime of C. Therefore (A + rbR)/A ⊆ Soc(R/A) and Soc(R/A) ≠ 0.

Conversely, assume that Soc(R/A) ≠ 0. Then there exists x ∈ R \ A such that xM ⊆ A for some maximal ideal M of R with A ⊆ M. Let B = A + xR. Then B/A ≅ R/M. Let C be an ideal of R containing A and maximal with respect to x ∉ C. Then C is completely irreducible and A = C ∩ B. Therefore C is a relevant completely irreducible divisor of A. □

Remark 1.12. The ring R of Example 1.7 is a zero-dimensional reduced ring for which Soc R = 0. Thus in this ring, the ideal (0) has no relevant completely irreducible divisors.

In Proposition 1.13, we observe a connection between relevant completely irreducible divisors of an ideal and maximal ideals that are strong Bourbaki associated primes of the ideal.

8 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

Proposition 1.13. Let A be a proper ideal of the ring R. A maximal ideal M of R is a strong Bourbaki associated prime of A if and only if there exists a relevant completely irreducible divisor of A that is M-primal.

Proof. Suppose C is a relevant completely irreducible divisor of A that is M-primal. The proof of Proposition 1.9 establishes the existence of an element u ∈ R \ A such that Mu ⊆ A. Hence M = A : u is a strong Bourbaki associated prime of A. Conversely, if M = A : x, then x ∉ A. Let C be an ideal of R that contains A and is maximal without x. Then M = C : x and C is M-primal. Since every ideal is an intersection of completely irreducible ideals, there exist completely irreducible ideals C_{i} such that A + Rx = ∩_{i∈I}C_{i}. Then (A + Rx)/A being simple implies that A = C ∩ (∩_{i∈I}C_{i}), where C is clearly a relevant component. □

2. Irredundant Intersections

We consider under what conditions an ideal may be represented as an irredundant intersection of completely irreducible ideals, where, as usual, irredundant means that none of the components may be omitted without changing the intersection.

If {B_{i}}_{i∈I} is a family of R-modules, then by an interdirect product of this family we mean an R-submodule of the direct product ∏_{i∈I} B_{i} that contains the direct sum ⨁_{i∈I} B_{i}. Similarly, if {R_{i}}_{i∈I} is a family of rings, by an interdirect product of this family of rings we mean a subring of the direct product ∏_{i∈I} R_{i} that contains the direct sum ⨁_{i∈I} R_{i} and the identity of the direct product.

In Lemma 2.1 and later in this section, we use the following notation. Let {C_{i}}_{i∈I} be a family of completely irreducible ideals of the ring R and let A = ⋂_{i∈I} C_{i}. For each j ∈ I let \overline{C}_{j} = ⋂_{i∈I,i≠j} C_{i}. We frequently use the following basic fact.

Lemma 2.1. Let A = ⋂_{i∈I} C_{i} be an irredundant representation of the proper ideal A with completely irreducible ideals C_{i}.

(i) There are elements u_{i} ∈ R \ C_{i} (i ∈ I) such that C_{i}^{∗} = C_{i} + Ru_{i} for i ∈ I and u_{i} ∈ C_{j} for all j ≠ i.

(ii) For each i ∈ I, the R-module \overline{C}_{i}/A has an essential socle generated by u_{i}+A. (iii) In the representation A = ⋂_{i∈I} C_{i}, no C_{i} can be replaced by a larger ideal

and still have the intersection be equal to A.

(iv) With u_{i} as in (i), let U_{i} = (Ru_{i} + A)/A. Then U_{i} is a simple R-module and the socle of R/A is an interdirect product of the U_{i}.

COMPLETELY IRREDUCIBLE IDEALS 9

Proof. (i) Since the intersection of the C_{i} (i ∈ I) is irredundant, for each i ∈ I there exists an element x_{i} ∈ \overline{C}_{i} such that x_{i} ∉ C_{i}. By Lemma 1.1, R/C_{i} is subdirectly irreducible, so some multiple rx_{i} ∉ C_{i} has the property that its image in R/C_{i} is in Soc (R/C_{i}). If we choose u_{i} = rx_{i} for each i ∈ I, then we obtain elements with the desired properties.

(ii) From (i) it is clear that ⟨u_{i} + A⟩ is in the socle of \overline{C}_{i}/A. By way of contradiction, suppose that there is a cyclic R-module ⟨v + A⟩ contained in \overline{C}_{i}/A and independent of ⟨u_{i} + A⟩. As C_{i} is completely irreducible, there is an r ∈ R such that rv − u_{i} ∈ C_{i}. Then rv − u_{i} ∈ C_{i} ∩ \overline{C}_{i} = A, contradicting the independence of ⟨v + A⟩ and ⟨u_{i} + A⟩.

(iii) If we replace C_{i} by a larger ideal, then the intersection will contain u_{i}, so it will no longer represent A.

(iv) As M_{i}u_{i} ⊆ A for a maximal ideal M_{i}, it is clear that the U_{i} = ⟨u_{i} + A⟩ are simple submodules of R/A. From (i) it follows that they are independent, so Soc (R/A) contains their direct sum D = ⊕_{i∈I}U_{i}. Suppose there is a v ∈ R such that v + A ∈ Soc (R/A), but not in D, say Mv ⊆ A for a maximal ideal M of R. Thus v ∉ C_{j} for some j ∈ I. If v ∉ C_{j} + Ru_{j}, then Mv ⊆ C_{j} implies C_{j} = (C_{j} + Ru_{j}) ∩ (C_{j} + Rv), contradicting the irreducibility of C_{j}. Hence v ∈ C_{j} + Ru_{j} for all j with v ∉ C_{j}. This means that the canonical injection R/A → ∏_{j∈I} R/C_{j} maps v to an element of ∏_{j∈I}(Ru_{j} + C_{j})/C_{j} ≅ ∏_{j∈I} U_{j}. □

We remark that part (iii) of Lemma 2.1 is also a consequence of the stronger statement: if A = B ∩ C, where C is irreducible and relevant for the decomposition of A, then C cannot be replaced by any proper overideal (see Noether [27, Hilfssatz II]). In fact, if C is properly contained in C′, then A = B ∩ C′ would lead to C = C + (B ∩ C′) = (C + B) ∩ C′, contradicting the irreducibility of C.

Remark 2.2. It will be useful to keep in mind that (iv) of Lemma 2.1 implies that

(1)  ⨁_{i∈I} Soc(\overline{C}_{i}/A) ⊆ Soc (R/A) ⊆ ∏_{i∈I} Soc(\overline{C}_{i}/A).

Notice also that Soc(\overline{C}_{i}/A) ≅ Soc(C_{i} + \overline{C}_{i})/C_{i} = Soc (R/C_{i}) = C_{i}^{∗}/C_{i}.

Next we exhibit an example where Soc(R/A) is the direct product of the Soc(\overline{C}_{i}/A).

Example 2.3. For each nonnegative integer n, let Z_{n} = ⟨x_{n}⟩ be a cyclic group of order p, where p is a fixed prime integer. Consider the product ∏_{n<ω} Z_{n} with trivial multiplication and let R be the ring obtained by adjoining an identity to ∏_{n<ω} Z_{n}

10 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

in such a way that as an additive group R = ℤ ⊕ ∏_{n<ω} Z_{n} and multiplication is defined as (a, b) × (c, d) = (ac, ad + cb), where a, c ∈ ℤ and b, d ∈ ∏_{n<ω} Z_{n}.

For each nonnegative integer m, let C_{m} ≅ ∏_{n<ω,n≠m} Z_{n} denote the subgroup of ∏_{n<ω} Z_{n} of elements having zero coordinate in component Z_{m}. Then C_{m} is an ideal of R that is maximal with respect to not containing the element that is x_{m} in component Z_{m} and zero in all other components. Thus C_{m} is completely irreducible. We have ⋂_{n<ω} C_{n} = 0 is an irredundant intersection, and in this example, the first containment relation in (1) is strict, while the second becomes equality.

For an irredundant intersection A = ⋂_{i∈I} C_{i} with completely irreducible ideals C_{i}, we demonstrate in Example 2.4 that Soc (R/A) may fail to be an essential submodule of R/A.

Example 2.4. There is a ring R in which the zero ideal is an irredundant intersection of completely irreducible ideals and Soc R is not an essential submodule of R. Let {p_{n} | n < ω} be the set of prime numbers, and let Z_{n} denote the ring ℤ/p_{n}^{2}ℤ (n < ω). In the direct product P = ∏_{n<ω} Z_{n} consider the subring R generated by 1 = (1, 1, …, 1, …), e = (p_{0}, p_{1}, …, p_{n}, …), and r_{n} = (0, …, 0, p_{n}, 0, …) for each n < ω. Then e^{2} = 0, er_{n} = 0 and r_{n}r_{m} = 0 for all n, m < ω. Consequently, a typical element of R has the form a = m·1+se+n_{0}r_{0} +⋯+n_{k}r_{k} for some k ≥ 0, where m, s, n_{i} ∈ ℤ. The ideal Re is just the infinite cyclic subgroup generated by e. Define C_{n} as an ideal containing all r_{i} (i < ω, i ≠ n), r_{n} − p_{0}^{n}p_{1}^{n} ⋯ p_{n}^{n}e, and being maximal with respect to r_{n} ∉ C_{n}. Clearly, the intersection ⋂_{n<ω} C_{n} contains no nonzero element of the socle Soc R = ⊕_{n<ω}Rr_{n}, and is evidently irredundant. To see that it cannot contain any element outside the socle either, assume by way of contradiction that a as above is contained in every C_{n}. Then either a multiple of se or ae = me is contained in every C_{n}, which is impossible, since p_{0}^{n}p_{1}^{n} ⋯ p_{n}^{n}e ∉ C_{n}. However, the socle of R is not essential in R, as the ideal Re intersects Soc R trivially.

Let A be a proper ideal of the ring R. In Theorem 2.5 we relate representations of A as an intersection of bigger ideals with representations of R/A as a subdirect product.

Theorem 2.5. For a proper ideal A of the ring R the following are equivalent: (i) A is the irredundant intersection of completely irreducible ideals;

COMPLETELY IRREDUCIBLE IDEALS 11

(ii) R/A is an irredundant subdirect product of subdirectly irreducible R-algebras R_{i}.

Proof. (i) ⇒ (ii) Let {C_{i}}_{i∈I} be a family of completely irreducible ideals of R such that A = ⋂_{i∈I} C_{i} is an irredundant intersection. Then R/A can be identified via the canonical map ϕ : R → ∏_{i∈I} R/C_{i} with the subring ϕ(R) of the direct product ∏_{i∈I} R/C_{i}, where each component R/C_{i} contains a simple R-module C_{i}^{∗}/C_{i} as an essential socle, so each component is subdirectly irreducible. If we drop a component R/C_{j}, then ϕ will no longer induce an embedding of R/A in the product ∏_{i∈I,i≠j} R/C_{i}, so the subdirect representation is irredundant.

(ii) ⇒ (i) Assume that R/A is an irredundant subdirect product of the subdirectly irreducible R-algebras R_{i} (i ∈ I). Let T = ∏_{i∈I} R_{i}, and let π_{i} : R/A → R_{i} be the ith canonical projection. The kernel of π_{i} is C_{i}/A, where C_{i} is a completely irreducible ideal of R properly containing A, and R_{i} is isomorphic to R/C_{i}. Clearly, A = ⋂_{i∈I} C_{i} and this representation is irredundant. □

Example 2.6. Let F_{p} denote the prime field of characteristic p, for a prime p. Define R as the direct product ∏_{p} F_{p} with p running over an infinite set of primes. Identify the product C_{q} of the F_{p} with p ≠ q (q a fixed prime) as a subset of R consisting of all tuples having a zero in position q. Under this identification, C_{q} is a completely irreducible ideal of R, and 0 = ⋂_{q} C_{q} is an irredundant intersection. (In this example, Soc R is an essential submodule of R.)

In the preceding example, we could have used equally well several subrings (containing ⊕_{p}F_{p}) of R, e.g. the von Neumann regular ring over which all von Neumann regular rings are unital algebras; see Fuchs-Halperin [13].

Another criterion for the existence of an irredundant intersection representation of an ideal A with completely irreducible ideals can be given in terms of the injective hull of R/A.

Theorem 2.7. The proper ideal A of the ring R admits an irredundant representation as an intersection of completely irreducible ideals C_{i} (i ∈ I) if and only if the injective hull E(R/A) of R/A is an interdirect product of injective hulls of simple R-modules.

Proof. The ‘only if’ part of the theorem is an immediate consequence of Theorem 2.5, since the components R/C_{i} in the subdirect product are R-modules each with a simple essential submodule.

12 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

Conversely, assume E(R/A) is an interdirect product of injective hulls E(S_{i}) of simple R-modules S_{i}. Define R_{i} = π_{i}(R/A), where π_{i} denotes the ith coordinate projection of the product of the E(S_{i}). Then R/A is evidently a subdirect product of the R_{i}, and the R_{i} are subdirectly irreducible. For each i, the submodule E(S_{i}) of E(R/A) has nonzero intersection with R/A, and this intersection is contained in the kernel of π_{j} for each j ≠ i. This shows that the representation is irredundant. By Theorem 2.5, A admits an irredundant representation as an intersection of completely irreducible ideals. □

For a proper ideal A of R, let S(A) denote the set of elements r ∈ R that are not prime to A, i.e. that satisfy A ⊂ A : r.

Proposition 2.8. Let A be a proper ideal of a ring R. If A is an irredundant intersection of completely irreducible ideals C_{i}, then S(A) is the union of maximal ideals of R, each of which is a strong Bourbaki associated prime of A.

Proof. Given r ∈ R, we have r ∈ S(A) ⟺ A ⊂ A : r. Since A : r = ⋂_{i∈I}(C_{i} : r), we have, by Lemma 2.1(iii), r ∈ S(A) ⟺ C_{i} ⊂ C_{i} : r for some i. Letting M_{i} = {t ∈ R | C_{i} ⊂ C_{i} : t}, the adjoint prime of C_{i}, we see that S(A) = ⋃_{i∈I} M_{i}.

The elements u_{i} ∈ R introduced in Lemma 2.1 satisfy M_{i} = C_{i} : u_{i} and u_{i} ∈ C_{j} for j ≠ i. Thus A : u_{i} = M_{i}, showing that M_{i} is a strong Bourbaki associated prime of the ideal A. □

We can now prove:

Corollary 2.9. Let A = ⋂_{i∈I} C_{i} be an irredundant intersection, where the C_{i} are completely irreducible ideals. Then A is an M-primal ideal if and only if all the C_{i} are M-primal.

Proof. If all the completely irreducible ideals C_{i} are M-primal, then by Proposition 2.8 S(A) = M, so A is M-primal. On the other hand, if A is M-primal, i.e. if S(A) = M, then again by Proposition 2.8, all the adjoint primes of the C_{i} must be equal to M. □

This leads to a coarser intersection decomposition of ideals admitting irredundant intersections with completely irreducible ideals.

Corollary 2.10. Assume that the proper ideal A of R has an irredundant representation as an intersection of completely irreducible ideals. Then A is an irredundant intersection of primal ideals with distinct adjoint maximal ideals such that each primal component is an irredundant intersection of completely irreducible ideals.

Proof. Let A = ⋂_{i∈I} C_{i} be an irredundant intersection, where the C_{i} are completely irreducible ideals. For each maximal ideal M that is the adjoint prime of some C_{i}, , let A^{[M]} denote the intersection of the C_{i} that are M-primal. By Corollary 2.9, A^{[M]} is M-primal. Thus A = ⋂_{M∈Max R} A^{[M]} is a decomposition of A into an irredundant intersection of primal ideals with distinct adjoint primes M, and each A^{[M]} is an irredundant intersection of completely irreducible ideals. □

Since the isolated M-component A_{(M)} of an ideal A is the intersection of all M-primal overideals of A, it is clear that A_{(M)} ⊆ A^{[M]} for each M.

It would be interesting to know if the converse of Corollary 2.10 is also true.

Example 2.11. It is possible that an ideal A is an irredundant intersection of completely irreducible ideals {C_{i}}_{i∈I} and also the intersection of relevant completely irreducible divisors {C′_{i}}_{i∈I′}, where no subset of {C′_{i}}_{i∈I′} gives an irredundant representation of A. Let R be the ring defined in Example 2.4. It has been proved there that its zero ideal admits a representation as an irredundant intersection of completely irreducible ideals. Define A_{n} (n > 0) to be an ideal containing all r_{i} (i < ω, i ≠ n), containing e, and maximal with respect to not containing r_{n}. Furthermore, for each n < ω let B_{n} (n < ω) be an ideal that contains all r_{i} with i > 0, contains the element r_{0} − p_{0}^{n}p_{1}^{n} ⋯ p_{n}^{n}e, and is maximal with respect to not containing r_{0}. In view of Proposition 1.9, all of A_{n}, B_{n} are relevant completely irreducible divisors of A. Evidently, A = ⋂_{n>0} A_{n} ∩ ⋂_{n<ω} B_{n}. Here each of the A_{n}, but none of B_{n}, is relevant for the intersection. Indeed, infinitely many B_{n} can be deleted from this intersection as long as infinitely many B_{n} remain. In view of this, this intersection cannot be made irredundant by canceling components.

3. The Question of Uniqueness

In her seminal paper [27] E. Noether proved that in a Noetherian ring every proper ideal A is the finite irredundant intersection of irreducible ideals, and this intersection has the following uniqueness properties: (1) the number of irreducible components is unique, (2) the components satisfy the Replacement Property as defined in (3.2), and (3) the components are primary ideals and their set of prime radicals (along with their multiplicities) is uniquely determined by A.

14 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

Our next goal is to prove, if possible, analogous results for ideals A that admit a representation as an irredundant intersection of completely irreducible ideals. Of course, the role of primary ideals will be taken over by primal ideals, and the set of prime radicals will be replaced by the set of adjoint maximal ideals.

In our attempt to generalize the Noetherian situation the first problem we face is that the cardinality of the set of completely irreducible components in an irredundant intersection need not be unique. This is illustrated by the following example.

Example 3.1. Let κ be an infinite cardinal and λ any cardinal such that κ < λ ≤ 2^{κ}. There exists a ring R that contains an ideal which is the irredundant intersection of κ completely irreducible ideals and also the irredundant intersection of λ completely irreducible ideals. Fix a prime integer p and let P = ∏_{α<κ} Z_{α} be the direct product of of κ cyclic groups Z_{α} of order p. Give P the trivial multiplication and make R = ℤ ⊕P into a ring as in Example 2.3. Evidently P = Soc R and |P | = 2^{κ}. For β < κ, the product P_{β} = ∏_{α<κ,α≠β} Z_{α} may be viewed as an ideal in R; it is completely irreducible, since it is maximal with respect to intersecting Z_{β} in zero. Clearly, 0 = ⋂_{β<κ} P_{β} is an irredundant intersection with κ components. Pick a subgroup P′ of P that contains the direct sum ⊕_{α<κ}Z_{α} and has index λ in P. Now P = P′ ⊕ ⨁_{γ<λ} W_{γ} for cyclic groups W_{γ} of order p. Consider completely irreducible ideals C_{α} that contain ⨁_{γ<λ} W_{γ} and are maximal with respect to intersecting Z_{α} in zero, and consider as well completely irreducible ideals C′_{γ} that contain P′ and are maximal with respect to intersecting W_{γ} in zero. Then the intersection 0 = ⋂_{α<κ} C_{α} ∩ ⋂_{γ<λ} C′_{γ} is irredundant with λ components.

As far as uniqueness of irredundant decompositions is concerned, we have the following:

Theorem 3.2. Let

A = ⋂_{i∈I} C_{i} = ⋂_{j∈J} B_{j}

be two irredundant intersection representations of the proper ideal A of the ring R with completely irreducible ideals C_{i}, B_{j}. Then:

(i) Each M ∈ Max R that occurs as the adjoint prime of some primal ideal C_{i} also occurs as the adjoint prime of some B_{j}. If it occurs a finite number of times in one intersection, then it occurs the same number of times in the other intersection.

COMPLETELY IRREDUCIBLE IDEALS 15

(ii) The intersections have the Replacement Property: for each C_{k} there is a B_{j} such that replacing C_{k} by B_{j} in the first intersection, we obtain an irredundant representation of A.

Proof. We start the proof with (ii). (Note that in the first paragraph of its proof irredundancy is irrelevant.)

(ii) Write \overline{C}_{k} = ⋂_{i≠k} C_{i} and A_{kj} = \overline{C}_{k} ∩ B_{j}. The R-module (C_{k} + \overline{C}_{k})/C_{k} ≅ \overline{C}_{k}/(C_{k} ∩\overline{C}_{k}) ≅ \overline{C}_{k}/A has an essential simple socle. Therefore, A = ⋂_{j∈J} A_{kj} with A_{kj} ⊆ \overline{C}_{k} implies that one of A_{kj} (j ∈ J) must be equal to A, say, \overline{C}_{k} ∩ B_{j} = A. Thus C_{k} can be replaced by B_{j}.

To show that if A = ⋂_{i∈I} C_{i} is irredundant, then so is the new intersection, assume by way of contradiction that C_{ℓ} (ℓ ≠ k) can be omitted from the new intersection, i.e. A = \overline{C}_{k} ∩ \overline{C}_{ℓ} ∩ B_{j} (where C_{k} and C_{ℓ} are missing). What has been proved is applied to this decomposition and the original one to conclude that here B_{j} can be replaced by one of the C_{i}. After we do this, at least one of C_{k} and C_{ℓ} would be missing from the intersection, contradicting the irredundancy of the decomposition A = ⋂_{i∈I} C_{i}.

(i) Both C_{k} and its replacement B_{j} satisfy \overline{C}_{k} ∩ C_{k} = A = \overline{C}_{k} ∩ B_{j}. Hence the covers satisfy C_{k}^{∗}/C_{k} ≅ Soc(\overline{C}_{k}/A) ≅ B_{j}^{∗}/B_{j}. It follows that both C_{k} and B_{j} are M-primal for the maximal ideal M that satisfies Soc(\overline{C}_{k}/A) ≅ R/M. Thus an M-primal completely irreducible ideal is always replaced by another M-primal ideal in an irredundant intersection.

Now if the first intersection contains n completely irreducible ideals that are M-primal, then successively replacing them by (necessarily distinct) completely irreducible ideals from the second intersection, it follows that the second intersection contains at least n M-primal ideals. Hence the claim is evident. □

Dilworth-Crawley [7, Theorem 4.2] prove an analogue of (ii) for compactly generated modular lattices under the hypothesis that the lattice is atomic (in our setting, this corresponds to the case where Soc (R/A) is essential in R/A). They also characterize atomic lattices where the components in irredundant intersections are unique. Since we do not have atomicity in general (see Example 2.4), in our search for uniqueness an independent approach is needed.

Remark 3.3. In [21] the rings are characterized for which every ideal can be represented uniquely as an irredundant intersection of completely irreducible ideals.

16 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

These rings are necessarily arithmetical. Indeed, it is shown in [21] that if A is the set of ideals of a ring R that are finite intersections of completely irreducible ideals, then R is arithmetical if and only if for each A ∈ A the representation of A as an irredundant intersection of completely irreducible ideals is unique.

4. When Every Ideal is an Irredundant Intersection

In this section we characterize the rings in which every ideal is an irredundant intersection of completely irreducible ideals. We start with the following observation.

Lemma 4.1. Consider the following statements about a proper ideal A of the ring R.

(i) A is an irredundant intersection of completely irreducible ideals.

(ii) Soc (R/A) is an essential submodule of R/A. Then (ii) implies (i). On the other hand, if (i) holds for every proper ideal of R, then also (ii) holds for every proper ideal of R.

Proof. Assume that (ii) holds and let E(R/A) denote the injective hull of R/A. Express Soc (R/A) as the direct sum ⊕_{i∈I}S_{i} of simple R-modules S_{i}, and let E(S_{i}) be a maximal essential extension of S_{i} in E(R/A). Then E(S_{i}) is an injective hull of S_{i} and ⊕_{i∈I}E(S_{i}) is an essential submodule of E(R/A). The canonical map of the direct sum ⊕_{i∈I}E(S_{i}) into the direct product ∏_{i∈I} E(S_{i}) extends to an injection of E(R/A) into ∏_{i∈I} E(S_{i}). Thus E(R/A) is an interdirect sum of the E(S_{i}). By Theorem 2.7, A is an irredundant intersection of completely irreducible ideals.

Assume that (i) holds for every proper ideal of R and let A be a proper ideal of R. By Lemma 2.1, Soc (R/A) cannot be trivial. Let B/A be a maximal essential extension of Soc (R/A) in R/A. Then R/B has trivial socle, so B cannot admit an irredundant representation as intersection of completely irreducible ideals. It follows that B = R. This means that Soc (R/A) is an essential submodule in R/A. □

Recall that a ring R is semi-artinian if every nonzero R-module contains a simple R-module. We will call a domain almost semi-artinian if every proper factor ring is semi-artinian.

Theorem 4.2. Every ideal of a ring is an irredundant intersection of completely irreducible ideals exactly if the ring is semi-artinian. Every nonzero ideal of an integral domain is an irredundant intersection of completely irreducible ideals if and only if the integral domain is almost semi-artinian.

Proof. Lemma 4.1 implies that all the ideals of a ring R have the stated property if and only if Soc (R/A) is essential in R/A for all proper ideals A of R. This is the case exactly if R is semi-artinian.

The claim concerning domains is an immediate consequence of the first part. □

Example 4.3. A ring that is perfect in the sense of Bass [2] is semi-artinian. Perfect rings can be defined in several equivalent ways; for example, the ring R is perfect is equivalent to each of the following statements: (i) every flat R-module is projective, (ii) R satisfies the minimum condition for principal ideals (see e.g. Anderson-Fuller [1, p. 315 ]).

Example 4.4. Consider the Bézout domain R of dimension 1 constructed by Heinzer-Ohm [20, Example 2.2]. For this ring R, Spec R is not Noetherian, but all of its localizations at maximal ideals are rank-one discrete valuation domains. Moreover, all but one of its maximal ideals are principal. It can be seen that this example is almost semi-artinian. Thus an almost semi-artinian domain need not have Noetherian spectrum.

More interesting examples are those almost semi-artinian domains that are also h-local, where an integral domain is h-local if each nonzero element is contained in only finitely many maximal ideals and each nonzero prime ideal is contained in a unique maximal ideal.

Example 4.5. The almost perfect domains introduced by Bazzoni-Salce [4] are defined as domains such that every proper factor ring is a perfect ring. They prove that an almost perfect domain can be characterized as an h-local domain R such that for every nonzero proper ideal A, the factor ring R/A contains a simple R-module. We are interested in obtaining more information about the ideal structure of almost perfect domains. An almost perfect domain is a Matlis domain of dimension ≤ 1 with Noetherian prime spectra, where an integral domain R is a Matlis domain if its field of fractions has projective dimension at most 1 as an Rmodule. The inclusion here is strict since a rank-one nondiscrete valuation domain is a Matlis domain of dimension 1 that fails to be almost perfect.

18 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

In an almost perfect domain, irreducible ideals are primary. Indeed, by [11] a necessary and sufficient condition for an irreducible ideal A of a ring R to be primary is that any strictly ascending chain of the form

A ⊂ A : b_{1} ⊂ A : b_{1}b_{2} ⊂ ⋯ ⊂ A : b_{1}b_{2} ⋯ b_{n} ⊂ …

for any sequence b_{1}, b_{2}, …, b_{n}, … in R terminates. This is evidently true if the descending chain condition holds for principal ideals mod A.

Since an almost perfect domain R is h-local, every nonzero ideal A of R is contained in at most a finite number of maximal ideals, M_{1}, …, M_{n}. Consequently, we can write

A = A_{(M_{1})} ∩ ⋯ ∩ A_{(M_{n})}.

The localization at a maximal ideal of a domain of dimension one is again of dimension one, and if R is a one-dimensional quasilocal domain with maximal ideal M, then every nonzero proper ideal of R is M-primary. Consequently, the isolated M_{i}-component A_{(M_{i})} must be M_{i}-primary for each i = 1, …, n.

We record as Theorem 4.6 an immediate consequence of [5, Theorem 8].

Theorem 4.6. In an almost perfect domain R, every nonzero ideal A is the product of pairwise comaximal primary ideals:

A = B_{1}B_{2} ⋯ B_{n},

where the B_{i} are primary ideals with distinct maximal ideals M_{i} as radicals. This product representation is unique up to the order of the factors. Since an almost perfect domain is almost semi-artinian, the B_{i} are irredundant intersections of completely irreducible M_{i}-primary ideals. □

In particular, it follows that an almost perfect domain is Laskerian. We recall that a ring is said to be strongly Laskerian if it is Laskerian and every primary ideal contains a power of its radical. If A is a nonzero ideal of a one-dimensional strongly Laskerian domain R, then R/A satisfies the minimum condition for principal ideals, for R/A ≅ ∏_{i=1}^{n} R_{i}, where each R_{i} is a zero-dimensional ring with nilpotent maximal ideal M_{i}. If r_{i} is a positive integer such that M_{i}^{r_{i}} = 0, then a strictly descending chain of principal ideals in R_{i} has length at most r_{i}. It follows that a strictly descending chain of principal ideals in R/A has length at most r_{1} + ⋯ + r_{n}. Thus R is almost perfect.

COMPLETELY IRREDUCIBLE IDEALS 19

Remark 4.7. The representations of the ideals B_{i} of Theorem 4.6 as irredundant intersections of completely irreducible M_{i}-primary ideals may be infinite intersections. For example, if R with maximal ideal M is a one-dimensional local strongly Laskerian domain that is not Noetherian, then M/M^{2} is infinite dimensional as a vector space over R/M. Since R is almost semi-artinian, every nonzero ideal of R is an irredundant intersection of completely irreducible ideals. However, every representation of M^{2} as an intersection of completely irreducible ideals is an infinite intersection (see for example part (iv) of Lemma 2.1).

5. Arithmetical Rings

Proposition 1.4 and Corollary 1.5 characterize the completely irreducible ideals of a Noetherian ring. In this section we give an explicit description of the completely irreducible ideals in a much different setting, that of an arithmetical ring. These rings arise naturally in the consideration of irreducible ideals. In Theorem 1.8 and Remark 1.6 of [14] it is shown that the following are equivalent: (i) the ring R is arithmetical, (ii) every primal ideal of R is irreducible, (iii) the ideal A_{(P)} is irreducible for every ideal A and prime ideal P of R with A ⊆ P.

We first consider the special case of a valuation ring, where by a valuation ring we mean a ring in which the ideals are linearly ordered with respect to inclusion.

Lemma 5.1. Let M be a maximal ideal of the ring R. The following statements are equivalent.

(i) R is a valuation ring with maximal ideal M.

(ii) The set of proper completely irreducible ideals of R is precisely the set of

ideals of the form rM, where r is a nonzero element of R. (iii) For every nonzero r ∈ R, the ideal rM is irreducible.

Proof. (i) ⇒ (ii) Let C be a completely irreducible ideal of R. Since C^{∗}/C is a simple R-module and R is a valuation ring, C^{∗} must be a principal ideal of R and C = MC^{∗}. On the other hand, every ideal of the form rM, r ∈ R and r nonzero, is completely irreducible by Theorem 1.3(v) and the fact that every ideal of a valuation ring is irreducible.

(ii) ⇒ (iii) This is clear.

(iii) ⇒ (i) We first observe that (iii) implies R is quasilocal. For if r is not in M but is in some other maximal ideal of R, then rM = M ∩ rR is an irredundant intersection, contradicting the fact that rM is irreducible.

20 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

Let x and y be nonzero elements of R. We show that the ideals xR and yR are comparable. Notice that M = xM : x and M = yM : y. The ideals xM and yM are irreducible, so by Theorem 1.3(v) xM and yM are completely irreducible. Thus it follows that xR and yR are the respective covers (unique minimal overideals) of xM and yM. Let I = (x, y)R and let A be an ideal of R with MI = Mx+My ⊂ A. Since xR is the cover of xM and yR is the cover of yM, we must have x ∈ A and y ∈ A. Also MI ⊂ I. Therefore I is the unique minimal overideal of MI, and MI is completely irreducible with cover I. In particular, MI is irreducible, so I/MI has dimension one as an R/M-vector space. Thus I = xR + MI or I = yR + MI. Since R is quasilocal, it follows that I = xR or I = yR, proving that xR and yR are comparable. □

Lemma 5.2. Let A be an ideal of a ring R. If M is a maximal ideal of R, then MA_{(M)} = (MA)_{(M)}.

Proof. Note first that MA_{(M)} ⊆ (MA)_{(M)}. For if m ∈ M and x ∈ A_{(M)}, then there exists y ∈ R\M such that xy ∈ A, so that mxy ∈ MA and mx ∈ (MA)_{(M)}. To complete the proof we verify locally that (MA)_{(M)} = MA_{(M)}. Clearly (MA)_{(M)}R_{M} = MAR_{M} = MA_{(M)}R_{M}, so suppose N is a maximal ideal of R distinct from M. Then MA_{(M)} ⊆ (MA)_{(M)} implies that A_{(M)}R_{N} = MA_{(M)}R_{N} ⊆ (MA)_{(M)}R_{N} ⊆ A_{(M)}R_{N}; hence MA_{(M)}R_{N} = (MA)_{(M)}R_{N}. □

Theorem 5.3. A ring R is arithmetical if and only if the completely irreducible proper ideals of R are precisely the ideals of the form MB_{(M)}, where M is a maximal ideal and B is a principal ideal having the property that BR_{M} ≠ 0.

Proof. Suppose that R is an arithmetical ring. Let C be a completely irreducible proper ideal of R and let M be the adjoint (maximal) ideal of C. By Theorem 1.3, CR_{M} is completely irreducible, and by Lemma 5.1, CR_{M} = (r/1)MR_{M}, where r ∈ R is such that its image r/1 in R_{M} is nonzero. Define B = rR. Since for every ideal A of R, the ideal A_{(M)} is the preimage of AR_{M} under the canonical mapping R → R_{M}, and since C = C_{(M)}, we have C = (MB)_{(M)}. Therefore by Lemma 5.2 C = MB_{(M)}, and every completely irreducible ideal of R has the stated form. By Lemmas 5.1 and 5.2 and Theorem 1.3 (vi), every ideal of the stated form is completely irreducible.

Conversely, assume that the completely irreducible proper ideals of R are precisely the ideals of the form MB_{(M)}, where M is a maximal ideal and B is a

COMPLETELY IRREDUCIBLE IDEALS 21

principal ideal of R having the property that BR_{M} ≠ 0. It suffices to prove that R_{M} is a valuation ring for every maximal ideal M of R. By Lemma 5.1 it suffices to show that (r/1)MR_{M} is irreducible for all elements r ∈ R such that r/1 ∈ R_{M} is nonzero. By assumption, M(rR)_{(M)} is an irreducible ideal of R, and this implies that (r/1)MR_{M} is an irreducible ideal of R_{M} (Remark 1.6 of [14]). □

An ideal A of a domain R is said to be archimedean if its only endomorphisms are multiplications by elements of R, i.e. if its endomorphism ring, End(A), is equal to R [18, page 71]. We observe in Remark 5.4 that completely irreducible ideals of a Prüfer domain need not be archimedean.

Remark 5.4. Let R = ℤ + xℚ[[x]] be the ring of power series in x with rational coefficients whose constant term is an integer. It is known that R is a Prüfer domain, dim R = 2 and P = xℚ[[x]] is the unique nonmaximal nonzero prime ideal of R. Moreover P is contained in each maximal ideal of R, and for p ∈ ℤ a prime integer, pR is a maximal ideal of R. Let V denote the valuation overring R_{2R} = ℤ_{2ℤ} + P, and consider the principal ideal A = xV of V. Since x ∈ P, A is also an ideal of R and is completely irreducible as an ideal of R with adjoint maximal ideal 2R and cover (x/2)V. Observe that End(A) is the proper valuation overring V of R. Therefore A is not archimedean as an ideal of R. Since A is completely irreducible, it is not possible to represent A as an intersection of archimedean ideals of R.

For a prime ideal P of a Prüfer domain, Fuchs and Mosteig prove in [17, Lemma 4.3] that the P -primal ideals form a semigroup under ideal multiplication. We generalize this result in Theorem 5.6. In the proof of Theorem 5.6, we use the following lemma.

Lemma 5.5. Let M be a maximal ideal of an arithmetical ring R. If A is a finitely generated ideal of R contained in M such that AR_{M} ≠ 0, then A_{(M)} is an M-primal ideal of R.

Proof. Observe that A_{(M)} is an M-primal ideal of R if and only if M is the unique maximal Krull associated prime of A_{(M)}. If P is a Krull associated prime of A_{(M)}, then necessarily P ⊆ M, so to show that A_{(M)} is an M-primal ideal it suffices to show that M is a Krull associated prime of A_{(M)}. Now by Lemma 2.4 of [14], M is a Krull associated prime of A_{(M)} if and only if MR_{M} is a Krull associated prime of AR_{M}. Also, a consequence of Lemma 1.5 in [15] is that every nonzero principal

22 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

ideal of a valuation ring V has adjoint prime the maximal ideal of V. Thus AR_{M} is MR_{M}-primal, and it follows that A_{(M)} is an M-primal ideal of R. □

Theorem 5.6. Let P be a prime ideal of an arithmetical ring R. If A and B are P -primal ideals such that for all maximal ideals M of R, ABR_{M} is a nonzero ideal of R_{M}, then AB is a P -primal ideal of R. In particular the set of regular P -primal ideals of R is a semigroup under ideal multiplication.

Proof. We first make two observations:

(a) If C is an ideal of R, then P is a Krull associated prime of C if and only if P R_{M} is a Krull associated prime of CR_{M} for some (equivalently, all) maximal ideal(s) M containing P. This is proved in Lemma 2.4 of [14].

(b) If C is a primal ideal of R, then P is the adjoint prime of C if and only if for each maximal ideal M of R containing P, CR_{M} is a P R_{M}-primal ideal of R_{M}. This is an application of (a) and the fact that an ideal C is P -primal if and only if P is the unique maximal member of the set of Krull associated primes of A.

We now prove the theorem. Suppose first that R is a valuation ring. Then AB is irreducible, hence primal, with adjoint prime, say Q. We claim that Q ⊆ P. If this is not the case then P ⊂ Q, and since AB is Q-primal, there exists q ∈ Q \ P such that qx ∈ AB for some x ∈ Q \ AB. Since B is P -primal, B : q = B. Also, B ⊆ qR, so B = qC for some ideal C. Necessarily B ⊆ C ⊆ B : q = B, so this forces B = C. Thus qB = B and qAB = AB. Since R is a valuation ring and x ∉ AB, we have AB ⊆ xR. Thus xqR ⊆ AB = qAB ⊆ xqR, and we conclude that xqR = AB. It follows that xq^{2}R = qAB = AB = xqR, so that xq = rxq^{2} for some r ∈ R. But this implies that (1 − rq)xq = 0, and since 1 − rq is a unit in R, xq = 0, contradicting the assumption that xqR = AB ≠ 0. Thus Q ⊆ P.

We show that in fact Q = P. If this is not the case, then Q ⊂ P and there exists p ∈ P \ Q. Since AB is Q-primal, AB : p^{2} = AB. Since A and B are P -primal, there exist x ∈ P \ A and y ∈ P \ B such that xp ∈ A and yp ∈ B. Hence xy ∈ AB : p^{2} = AB. Since R is a valuation ring, A ⊂ xR and B ⊂ yR. We conclude that xyR ⊆ AB ⊆ xyR. Thus by Lemma 5.5 and the fact that AB ≠ 0, we have that AB is an M-primal ideal of R. This forces Q = P = M, a contradiction to the assumption that Q ⊂ P. Hence P = Q and AB is P -primal.

We consider now the general case in which R is not necessarily a valuation ring. As noted above in (b), to prove that AB is P -primal, it suffices to show that P is the unique maximal member of the set of Krull associated primes of AB. By

COMPLETELY IRREDUCIBLE IDEALS 23

(b) AR_{M} and BR_{M} are P R_{M}-primal ideals of R, and since R_{M} is a valuation ring we have by the case considered above that ABR_{M} is a P R_{M}-primal ideal. By (a) P is a Krull associated prime of AB. To complete the proof it remains to show that P is the unique maximal member of the set of Krull associated primes of AB. Suppose that Q is a Krull associated prime of AB, and let M be a maximal ideal of R containing Q. By (a) QR_{M} is a Krull associated prime of ABR_{M}. Since ABR_{M} is P R_{M}-primal, QR_{M} ⊆ P R_{M}, so that Q ⊆ P. This completes the proof that every Krull associated prime of AB is contained in P. □

Remark 5.7. Simple examples show that even in a valuation ring R with maximal ideal M it is possible to have a proper ideal A that is completely irreducible (and therefore M-primal) such that A^{2} is not M-primal. Let V be a valuation domain whose value group is Γ = ℤ ⊕ ℤ with lexicographic ordering. Then the maximal ideal of V is principal generated by an element x where the value v(x) = (0, 1) is the smallest positive element in Γ. Let y ∈ V have value (1, 0), and let Q be the prime ideal of V of height one. Thus Q is generated by elements z having value v(z) = (1, b), where b ∈ ℤ may be negative. Notice that Q^{2} ⊂ yV ⊂ Q and that Q^{2} is Q-primary. Define R = V/Q^{2}. Then R is a valuation ring with maximal ideal M = xR. The ideal A = yR is completely irreducible, so M-primal. However, A^{2} = (0) is not M-primal. For a specific realization of this example, let x, y be indeterminates over a field k, let W = k(x)[y]_{(y)}, let Q = yW, and set V = k[x]_{(x)} + Q. Then V is a rank-two valuation domain with principal maximal ideal xV and height-one prime Q. The valuation ring R = V/Q^{2} has maximal ideal M = xR. Since Q^{2} is Q-primary, the ideal (0) of R is not M-primal. The principal ideal A = yR is a completely irreducible ideal and therefore is M-primal. However, A^{2} = (0) is not M-primal.

If we consider only those completely irreducible regular ideals whose adjoint prime is a fixed maximal ideal M, then we can verify:

Theorem 5.8. Let R be an arithmetical ring and let M be a maximal ideal of R. The set F of completely irreducible regular ideals of R with adjoint maximal ideal M is closed under ideal-theoretic multiplication, and F with this multiplication is a totally ordered cancellative semigroup.

Proof. We show first that: (⋆) If I and J are principal ideals of R and I_{(M)} and J_{(M)} are regular ideals of R, then (IJ)_{(M)} = I_{(M)}J_{(M)}. Suppose that I or J, say

24 LASZLO FUCHS, WILLIAM HEINZER, AND BRUCE OLBERDING

I, is not contained in M. Then IJR_{M} = JR_{M}, so that (IJ)_{(M)} = J_{(M)}, and since I ⊈ M, I_{(M)} = R. Thus if I or J is not contained in M, then (⋆) holds. Otherwise, both I and J are contained in M, and by Lemma 5.5 I_{(M)} and J_{(M)} are M-primal ideals of R. As a product of regular M-primal ideals, I_{(M)}J_{(M)} is M-primal by Theorem 5.6. Thus (I_{(M)}J_{(M)})_{(M)} = I_{(M)}J_{(M)}, so since IJ ⊆ I_{(M)}J_{(M)} we have that (IJ)_{(M)} ⊆ I_{(M)}J_{(M)}, and it remains to verify the reverse inclusion. Suppose that x ∈ I : b and y ∈ J : c for some b, c ∈ R \ M. Then xy ∈ IJ : bc ⊆ (IJ)_{(M)} since bc ∉ M. It follows then that I_{(M)}J_{(M)} ⊆ (IJ)_{(M)}, and this proves (⋆).

Now let A and B be in F. Then AB is an M-primal ideal by Theorem 5.6. By Theorem 5.3, A = MJ_{(M)} and B = MK_{(M)} for some principal ideals J and K of R. Since A and B are regular ideals, so are J_{(M)} and K_{(M)}. Now if M ≠ M^{2}, then since by Theorem 5.6, M^{2} is primal, hence irreducible, it follows that M/M^{2} is a onedimensional R/M-vector space. Hence if M ≠ M^{2}, M = I_{(M)} for some principal ideal I ⊆ M. In this case we have by (⋆) (and the fact that M is regular) that AB = I_{(M)}J_{(M)}K_{(M)}M = (IJK)_{(M)}M, so AB ∈ F by Theorem 5.3. Otherwise, if M = M^{2}, then again by (⋆) AB = (IJ)_{(M)}M and by Theorem 5.3, AB ∈ F.

Now assume AC = BC, where A, B, C are in F. Using Theorem 5.3, write A = (MJ)_{(M)}, B = (MK)_{(M)} and C = (ML)_{(M)} for principal ideals J, K, and L of R. Then ACR_{M} = BCR_{M} implies that M^{2}JLR_{M} = M^{2}KLR_{M}. Moreover LR_{M} is a regular principal ideal of R_{M} since C is regular and C ⊆ L_{(M)} implies CR_{M} ⊆ LR_{M}. Thus LR_{M} is a cancellation ideal of R_{M} and JM^{2}R_{M} = KM^{2}R_{M}. If MR_{M} = M^{2}R_{M}, then clearly AR_{M} = CR_{M}. Otherwise, if MR_{M} ≠ M^{2}R_{M}, then the irreduciblity of M^{2}R_{M} implies (as above) that MR_{M} is principal and regular, in which case also AR_{M} = CR_{M}. Therefore A = A_{(M)} = C_{(M)} = C.

Since R is an arithmetical ring, the ideals of R_{M} are linearly ordered with respect to inclusion and thus so are their preimages under the canonical map R → R_{M}. Therefore for ideals A and B of R, the ideals A_{(M)} and B_{(M)} of R are comparable. It follows that the ideals in F are totally ordered under inclusion. □

Acknowledgement. We thank the referee for numerous helpful suggestions about this paper.

References

[1] F.W. Anderson and K.R. Fuller, Rings and Categories of Modules, Springer, 1974.

[2] H. Bass, Finitistic dimension and a homological generalization of semiprimary rings, Trans. Amer. Math. Soc. 95 (1960), 466-488.

[3] S. Bazzoni, Divisorial domains, Forum Math. 12 (2000), 397-419.

[4] S. Bazzoni and L. Salce, Almost perfect domains, Coll. Math. 95 (2003), 285-301.

[5] J. Brewer and W. Heinzer, On decomposing ideals into products of comaximal ideals, Comm. Algebra 30 (2002), 5999-6010.

[6] R.P. Dilworth, Abstract commutative ideal theory, Pacific J. Math. 12 (1962), 481-498.

[7] R.P. Dilworth and P. Crawley, Decomposition theory for lattices without chain conditions, Trans. Amer. Math. Soc. 96 (1960), 1-22.

[8] M. Fontana and E. Houston, On integral domains whose overrings are Kaplansky ideal transforms, J. Pure Appl. Algebra 163 (2001), 173-192.

[9] M. Fontana, J.A. Huckaba, I.J. Papick and M. Roitman, Prüfer domains and endomorphism rings of their ideals, J. Algebra 157 (1993), 489-516.

[10] M. Fontana and S. Kabbaj, Essential domains and two conjectures in ideal theory, to appear Proc. Amer. Math. Soc.

[11] L. Fuchs, A condition under which an irreducible ideal is primary, Quart. J. Math. Oxford 19 (1948), 235-237.

[12] L. Fuchs, On primal ideals, Proc. Amer. Math. Soc. 1 (1950), 1-6.

[13] L. Fuchs and I. Halperin, On the imbedding of a regular ring in a regular ring with identity, Fund. Math. 54 (1964), 285-290.

[14] L. Fuchs, W. Heinzer and B. Olberding, Commutative ideal theory without finiteness conditions: primal ideals, to appear Trans. Amer. Math. Soc.

[15] L. Fuchs, W. Heinzer and B. Olberding, Maximal prime divisors in arithmetical rings, in Rings, Modules, Algebras, and Abelian Groups, Proceedings of the Algebra Conference - Venezia, Marcel Dekker vol.236 pages 189-203.

[16] L. Fuchs, W. Heinzer and B. Olberding, Commutative ideal theory without finiteness conditions: irreducibility in the quotient field, in preparation.

[17] L. Fuchs and E. Mosteig, Ideal theory in Prüfer domains — an unconventional approach, J. Algebra 252 (2002), 411-430.

[18] L. Fuchs and L. Salce, Modules over non-Noetherian Domains, Math. Surveys 84 (2001).

[19] S. Gabelli and E. Houston, Ideal theory in pullbacks, Non-Noetherian commutative ring theory, 199-227, Math. Appl. 520, Kluwer Acad. Publ. Dordrecht 2000.

[20] W. Heinzer and J. Ohm, Locally Noetherian commutative rings, Trans. Amer. Math. Soc. 158 (1971), 273-284.

[21] W. Heinzer and B. Olberding, Unique irredundant intersections of completely irreducible ideals, in preparation.

[22] J. Iroz and D. Rush, Associated prime ideals in non-Noetherian rings, Can. J. Math. 36 (1984), 344-360.

[23] W. Krull, Idealtheorie in Ringen ohne Endlichkeitsbedingung, Math. Ann. 101 (1929), 729-744.

[24] W. Krull, Idealtheorie, Ergebnisse d. Math. (Berlin, 1935).

[25] A. Loper and M. Roitman, The content of a Gaussian polynomial is invertible, preprint.

[26] T. Lucas, Invertibility of the content of a Gaussian polynomial, preprint.

[27] E. Noether, Idealtheorie in Ringbereichen, Math. Ann. 83 (1921), 24-66.

Department of Mathematics, Tulane University, New Orleans, Louisiana 70118

E-mail address: fuchs@tulane.edu

Department of Mathematics, Purdue University, West Lafayette, Indiana 47907

E-mail address: heinzer@math.purdue.edu

Department of Mathematical Sciences, New Mexico State University, Las Cruces, New Mexico 88003-8001

E-mail address: olberdin@emmy.nmsu.edu